3x-4y=12 and x=2y-4
Answer by angphy(3)

Substitute in the value for x.
3(2y-4)-4y=12
6y-12-4y=12 Then, combine like terms
2y-12=12 Add 12 to both side of the equal sign to isoslate the variable y
+12 +12
2y=24 Divide by 2 to get the value of y
y=12. Substitute this value into the 2nd equation
x=2(12)-4
x=24-4
x=20.
Put it in ordered pairs for your answer
(20,12)